How long does it take a child on a swing to complete one swing if her center of gravity is 4.00 m



below the pivot? Explain your answer. (

1
Expert's answer
2021-12-23T10:50:03-0500

We can find the period of one swing of a child from the definition of the period of pendulum:


"T=2\pi\sqrt{\dfrac{l}{g}}=2\pi\sqrt{\dfrac{4\ m}{9.8\ \dfrac{m}{s^2}}}=4.01\ s."